# justification=centering

### v1

@@ 5 Noise behavior in RSM
In HCI, detection limits are commonly set at a $5\sigma$ level, corresponding to a false alarm probability (FAP) of $3\times 10^{-7}$ based on a Gaussian noise assumption in processed frames (2014ApJ...792...97M). However, this approach may not accurately reflect the true noise characteristics in HCI datasets. Alternative approaches, based e.g. on Laplacian distributions (2019MNRAS.487.2262P; 2020A&A...633A..95D), have been proposed to address this limitation, especially at small separations. Moreover, methods like those in 2018AJ....155...19J, 2023AJ....166...71B, and 2024A&A...692A.126D avoid assuming any specific noise model, instead using adaptive techniques to derive thresholds directly from data, reducing the reliance on potentially mismatched distributions.
In the context of RSM, fitting a theoretical distribution to the noise in the final RSM maps (hereafter referred to as RSM noise) is challenging due to the nonlinear dependency of the final RSM score on the likelihood functions and transition probabilities, leading to the distortion of any pre-assumed distribution. 2021A&A...646A..49D defines the detection threshold for the contrast curve as the brightest false positive (FP) in the final detection map. As noted in Appendix F, this threshold is broadly consistent with those defined in this section. However, it lacks robustness in the presence of astrophysical signals and is not designed to detect them. An alternative approach explored by 2022A&A... [^6] A..33D , based on reversed parallactic angles, was also considered as a potential threshold applicable to both detection maps and contrast curves. However, it posed challenges by revealing a high number of false positives, ultimately limiting its utility as a reliable detection criterion.
@@
5.1 Fitting a log-normal distribution to the RSM noise
In this method, we take advantage of the large amount of data in the SHINE F150 survey to build distribution samples at each angular separation. Following 2023AJ....166...71B, we treat independent pixel realizations, i.e., the central pixels of non-overlapping resolution elements, as distribution samples. Noise analysis at a given separation across the entire survey shows multiple RSM noise distributions, leading us to analyze realizations per cluster and per combination of PSF subtraction techniques in RSM, at each considered angular separation.
The final RSM maps contain only positive values, leading to a noise distribution that is not centered on zero and exhibits positive skewness. After testing multiple distributions as shown in Appendix G, we selected the log-normal distribution for its minimal parameter requirements and its consistency with the shape of the RSM noise distribution across separations and clusters. To define a detection threshold, we set a $3\times 10^{-7}$ FAP for the log-normal distribution, aligning with the conventional $5\sigma$ threshold in Gaussian statistics. To prevent bias when fitting the distribution, we exclude scores above 0.1 from the distribution samples, as these correspond to bright signals.
Figure 3: Quantile-Quantile Plot of the RSM noise in Cluster 1 at $0\aas@@fstack{\prime\prime}25$ (left) and $1\arcsec$ (right) using APCA. The plot compares the RSM noise distribution to the Lognormal Distribution, excluding values above the $3\times 10^{-7}$ False Alarm Probability of the fit. The Coefficient of Determination ($R^{2}$) is also displayed, indicating a good fit.
@@ 5.1 Fitting a log-normal distribution to the RSM noise
Figure 4: Top. Log-normal distributions obtained by fitting RSM noise histograms at $0\aas@@fstack{\prime\prime}25$ for various observing conditions (clusters) using RSM with APCA. Bottom. Detection thresholds derived from the lognormal distributions at $3\times 10^{-7}$ FAP, as a function of angular separation and of observing conditions.
In Figure 4, we show how varying observing conditions impact the RSM noise distribution and, consequently, the chosen detection threshold. Figure 4(a) presents the log-normal fits for different clusters at $0\aas@@fstack{\prime\prime}25$ using RSM with APCA. Cluster 3, featuring a large number of frames per target, shows a narrower distribution and noise tail, while cluster 4, with mostly poor-quality observations, shows a broader tail and wider distribution, indicating a higher detection threshold. This plot highlights the varying RSM noise behaviors across observations under different conditions. Figure 4(b) illustrates the variations of the $3\times 10^{-7}$ FAP threshold across clusters and separations, highlighting the advantage of setting separate thresholds per separation to accommodate distinct noise properties. It also emphasizes the benefit of adjusting thresholds for observations under varying conditions. This approach, consistent with predictions, results in lower thresholds for Cluster 3 and higher ones for Cluster 4 and 5. This analysis underscores the importance of considering observing conditions when setting thresholds, as applying a single threshold across all datasets could lead to significant under- or over-estimations of the detection thresholds, as evidenced by the almost two orders of magnitude difference in thresholds between the best and worst clusters.
@@ 5.2 Threshold using maximum F1 score
In this section, we explore the possibility of defining a detection threshold without assuming a specific RSM noise distribution, an approach particularly relevant for samples containing a limited number of observations, where robust noise histograms cannot be built, especially for small angular separations. Following 2024A&A...692A.126D, we define detection thresholds by empirically maximizing the F1 score at each angular separation through fake companion injections. The F1 score is defined as follows:
@@
Figure 5: Variation of the true positive rate (TPR), false positive rate (FPR), and F1 score for the cluster 1 center at $0\aas@@fstack{\prime\prime}7$. The black dashed line marks the threshold that maximizes the fitted F1 score curve (grey curve), balancing high true positive (blue curve) recovery with minimal false positives (red curve).
To compute the false positives, all independent pixel realizations at each separation are considered as devoid of circumstellar signal (pure noise). Varying the threshold, we calculate the number of aperture values exceeding the threshold, which represent the number of false positives in the considered annulus. The false positive rate (FPR) is then defined as the ratio of these false positives to the total number of independent realizations. For true positives, we follow 2014ApJ...792...97M by first estimating noise levels in the residual cube produced by APCA. We then inject synthetic point sources with fluxes between 2 and 3 $\sigma$ of the measured noise. An injected source is considered recovered if its RSM score surpasses the considered detection threshold, and the number of such recoveries within the annulus defines the true positives. The optimal threshold is determined by maximizing the F1 score, ensuring a balance between high true positive recovery and minimal false positives.
Figure 5 shows the evolution of TPR, FPR, and the F1 score for varying thresholds, applied to the cluster 1 center (corresponding to the BD-15 705 dataset) using RSM with APCA at $0\aas@@fstack{\prime\prime}7$ separation. The figure highlights the fact that the optimal F1 score favors minimal FPR and high TPR, thereby effectively distinguishing noise from potential true positives. This metric is highly dataset-specific as it relies heavily on the false positive rate and the true positive rate, the latter also depending on the flux of injected signals, which is influenced by the noise distribution. Consequently, this approach captures the unique characteristics of each dataset and aligns the threshold closely with the specific noise properties of the data. This technique has however several limitations. First, while it minimizes the false positive rate, it does not explicitly control the acceptable number of false positives per annulus. Additionally, it assumes all independent realizations are false positives, which may not hold if a real signal is present. In such cases, the maximum F1 score metric tends to exclude the true signal by minimizing false positives. One solution is to remove bright signals by injecting negative fake companions to avoid bias. Finally, this approach requires noise computation, fake companion injections, and a complete RSM processing for each separation, making it computationally expensive.
@@ 5.3 Comparison and reliability of the detection thresholds
Figure 6: Comparison of detection thresholds with false positives and injected companions for the cluster 1 center at an angular separation of $0\aas@@fstack{\prime\prime}6$, showing the RSM scores of injected fake companions (gold dots), false positives (red crosses), the maximum F1 score threshold (dashed blue), and the lognormal threshold (solid blue).
@@
While it arguably provides the most relevant thresholds for a given data set, the maximum F1 score is computationally demanding to apply on each dataset of the survey, and too specific to generalize across an entire cluster (see more about that in Appendix H). Conversely, the detection thresholds derived from a log-normal distribution are more conservative, computationally efficient, and reflect the noise behavior under similar observing conditions. However, they rely on a predefined distribution, large sample sizes, and do not capture dataset-specific nuances. For the present study, focusing on a large survey, we adopt the threshold derived from the $3\times 10^{-7}$ FAP of the log-normal distribution, and derive it at each separation, for each combination of PSF subtraction technique, and for each cluster. We recommend using the F1 score threshold for smaller samples, while noting its limitations described in Sect. 5.2.
@@ 6 Sensitivity limits
In this study, we adopt the 50% completeness curve described in 2021A&A...646A..49D, also known as contrast curves, to assess the sensitivity limits to point-like sources. Such curves indicate the flux at which 50% of injected signals are detected above the selected threshold, here the $3\times 10^{-7}$ FAP from the log-normal distribution, which corresponds to the $5\sigma$ contrast curve when assuming Gaussian noise. Here, we analyze the effects of threshold selection and PSF subtraction technique combinations on contrast curves, define the optimal RSM contrast curve, and evaluate its reliability.
@@ 6.1 Influence of threshold and PSF subtraction techniques
Figure 8: Contrast comparison at 50% completeness using thresholds based on the maximum F1 score (blue), the log-normal RSM noise assumption (orange), and the $5\sigma$ APCA contrast curve (black) across different PSF subtraction techniques applied to the center of cluster 1.
@@
To evaluate how the chosen threshold impacts contrast curves, Fig. 8 presents a comparison for the center of cluster 1. It includes contrast curves derived using the maximum F1 score threshold (blue) and the lognormal threshold (orange) across four combinations of PSF subtraction techniques: APCA, APCA-NMF, APCA-NMF-LOCI, and APCA-NMF-LOCI-FMKLIP-FMLOCI. A standard $5\sigma$ contrast curve based on full-frame PCA with five principal components is shown in black for comparison. The RSM contrast curves show a good level of consistency between the two ways to define the detection threshold. This is especially the case when multi-technique combinations are used (see bottom panels in Fig. 8), while the log-normal threshold leads to slightly worse sensitivity than the F1-score one when RSM is performed with APCA only (Fig. 8 top left) due to the higher threshold in that case. Consistency was also found between the log-normal contrast curves derived here and the contrast curves based on the appearance of a first false positive proposed in 2022A&A... [^6] A..33D for the SHARDDS survey (see Appendix F).
@@
Figure 9: Effect of different combinations of PSF subtraction techniques in RSM on the fitted lognormal distributions, detection thresholds, and contrast curves. Top left shows the difference of lognormal fit across different PSF subtraction techniques with colors ranging from dark violet for multi combinations to yellow for singular combinations. Top right shows the $3\times 10^{-7}$ FAP under lognormal distribution for the different combinations across all separations. Bottom left shows the contrast curve delivered by the different combinations with 50% completeness.
@@
The choice of which PSF subtraction techniques to combine is an important aspect of the RSM algorithm, which is not straightforward to optimize. Instead of relying on a single combination identified as optimal by the RSM framework, we propose to systematically evaluate the performance of various combinations of PSF subtraction techniques. We analyze the differences in noise distributions in the final frames, the corresponding thresholds, and the contrast curves achievable with each combination. Figure 9 examines these effects across the 28 possible combinations of PSF subtraction techniques, ranging from single PSF subtraction techniques (in yellow) to multi-combinations (in dark violet). Figure 9 (top left) displays the median lognormal fits for each combination, clearly showing narrower noise distributions for multi-combinations and broader distributions for single techniques, indicating improved performance with combined techniques. Figure 9 (top right) presents the thresholds derived from these fits across all combinations and separations. Noise thresholds, on average, range from approximately $10^{-4}$ for multi-combinations to nearly 10 for LOCI – exceeding RSM’s maximum score. This renders LOCI non-physical and unreliable when used alone, a consequence of its high false positive rate in RSM for many datasets. Figure 9 (bottom left) illustrates the contrast curves for the various combinations, showing an improvement by a factor of 3 for multi-combinations. Contrast values with lognormal thresholds exceeding 0.1 in RSM were excluded from the bottom-left panel due to difficulties in achieving convergence. This further demonstrates that incorporating multiple PSF subtraction techniques in RSM significantly enhances detection limits. Moreover, it highlights RSM’s ability to leverage the diverse noise patterns of different PSF subtraction methods, effectively reducing the misclassification of transient speckles as planetary signals. RSM’s efficiency in integrating additional information – either by combining different PSF subtraction techniques or by increasing the number of frames – is evident through narrower noise distributions, improved thresholds, and enhanced detection limits (see also Figure 4(a)).
@@ 6.2 Optimal usage of RSM
Figure 11: Optimal contrast curve achieved by different combinations of PSF subtraction techniques in RSM. Grey curves show the median contrast performance for each combination, computed across all datasets in cluster 1. The red curve highlights the best achievable performance at each separation, given by the median of the minimum contrast values across all combinations. The black curve shows the median of the maximum contrast values delivered by the two best-performing combinations: APCA-NMF-LOCI-FMKLIP and APCA-LOCI-FMKLIP.
@@
Figure 12: Comparison of recovered fake point sources in cluster center 1 using RSM maps with different PSF subtraction technique combinations: (a) APCA-NMF-LOCI-FMKLIP and (b) APCA-LOCI-FMKLIP. These are compared to the signal-to-noise ratio PCA map with five principal components (c). The RSM maps shown in this figure are thresholded (RSM map minus the separation-dependent threshold). Injected signals recovered above the detection threshold are marked with yellow squares, while non-recovered signals appear in red squares.
Figure 13: Injected fake point sources compared to the optimal RSM contrast curve (black) and PCA 5-component contrast curve (red) for cluster center 1

### v2

@@ 5 Noise behavior in RSM
In HCI, detection limits are commonly set at a $5\sigma$ level, corresponding to a false alarm probability (FAP) of $3\times 10^{-7}$ based on a Gaussian noise assumption in processed frames (Mawet et al., 2014). However, this approach may not accurately reflect the true noise characteristics in HCI datasets. Alternative approaches based on, for example, Laplacian distributions (Pairet et al., 2019; Dahlqvist et al., 2020) have been proposed to address this limitation, especially at small separations. Moreover, methods such as those in Jensen-Clem et al. (2018), Bonse et al. (2023), and Daglayan et al. (2024) avoid assuming any specific noise model, instead using adaptive techniques to derive thresholds directly from data, reducing the reliance on potentially mismatched distributions.
In the context of RSM, fitting a theoretical distribution to the noise in the final RSM maps (hereafter referred to as RSM noise) is challenging due to the nonlinear dependency of the final RSM score on the likelihood functions and transition probabilities, leading to the distortion of any pre-assumed distribution. Dahlqvist et al. (2021b) defines the detection threshold for the contrast curve as the brightest false positive (FP) in the final detection map. As noted in Appendix E, this threshold is broadly consistent with those defined in this section. However, it lacks robustness in the presence of astrophysical signals and is not designed to detect them. An alternative approach explored by Dahlqvist et al. (2022) , based on reversed parallactic angles, was also considered as a potential threshold applicable to both detection maps and contrast curves. However, it posed challenges by revealing a high number of false positives, ultimately limiting its utility as a reliable detection criterion.
@@
5.1 Fitting a lognormal distribution to the RSM noise
In this method, we take advantage of the large amount of data in the SHINE F150 survey to build distribution samples at each angular separation. Following Bonse et al. (2023), we treat independent pixel realizations, i.e., the central pixels of non-overlapping resolution elements, as distribution samples. Noise analysis at a given separation across the entire survey shows multiple RSM noise distributions, leading us to analyze realizations per cluster and per combination of PSF subtraction techniques in RSM, at each considered angular separation.
The final RSM maps contain only positive values, leading to a noise distribution that is not centered on zero and exhibits positive skewness. After testing multiple distributions as shown in Appendix F, we selected the lognormal distribution for its minimal parameter requirements and its consistency with the shape of the RSM noise distribution across separations and clusters. To define a detection threshold, we set a $3\times 10^{-7}$ FAP for the lognormal distribution, aligning with the conventional $5\sigma$ threshold in Gaussian statistics. To prevent bias when fitting the distribution, we exclude scores above 0.1 from the distribution samples, as these correspond to bright signals.
Figure 3: Quantile-quantile plot of the RSM noise in Cluster 1 at $0\aas@@fstack{\prime\prime}25$ (left) and $1\arcsec$ (right) using APCA. The plot compares the RSM noise distribution to the lognormal distribution, excluding values above the $3\times 10^{-7}$ FAP of the fit. The coefficient of determination ($R^{2}$) is also displayed, indicating a good fit.
@@ 5.1 Fitting a lognormal distribution to the RSM noise
Figure 4: Top. Lognormal distributions obtained by fitting RSM noise histograms at $0\aas@@fstack{\prime\prime}25$ for various observing conditions (clusters) using RSM with APCA. Bottom. Detection thresholds derived from the lognormal distributions at $3\times 10^{-7}$ FAP, as a function of angular separation and of observing conditions.
In Figure 4, we show how varying observing conditions impact the RSM noise distribution and, consequently, the chosen detection threshold. Figure 4(a) presents the lognormal fits for different clusters at $0\aas@@fstack{\prime\prime}25$ using RSM with APCA. Cluster 3, featuring a large number of frames per target, shows a narrower distribution and noise tail, while cluster 4, with mostly poor-quality observations, shows a broader tail and wider distribution, indicating a higher detection threshold. This plot highlights the varying RSM noise behaviors across observations under different conditions. Figure 4(b) illustrates the variations of the $3\times 10^{-7}$ FAP threshold across clusters and separations, highlighting the advantage of setting separate thresholds per separation to accommodate distinct noise properties. It also emphasizes the benefit of adjusting thresholds for observations under varying conditions. This approach, consistent with predictions, results in lower thresholds for Cluster 3 and higher ones for Cluster 4 and 5. This analysis underscores the importance of considering observing conditions when setting thresholds, as applying a single threshold across all datasets could lead to significant underestimations or overestimations of the detection thresholds, as evidenced by the almost two orders of magnitude difference in thresholds between the best and worst clusters.
@@ 5.2 Threshold using maximum F1 score
In this section, we explore the possibility of defining a detection threshold without assuming a specific RSM noise distribution, an approach particularly relevant for samples containing a limited number of observations, where robust noise histograms cannot be built, especially for small angular separations. Following Daglayan et al. (2024), we define detection thresholds by empirically maximizing the F1 score at each angular separation through fake companion injections. The F1 score is defined as follows:
@@
Figure 5: Variation of the TPR, FPR, and F1 score for the cluster 1 center at $0\aas@@fstack{\prime\prime}7$. The dashed purple line marks the threshold that maximizes the fitted F1 score curve (dark blue curve), balancing high true positive (orange curve) recovery with minimal false positives (dark red curve).
To compute the false positives, all independent pixel realizations at each separation are considered as devoid of circumstellar signal (pure noise). Varying the threshold, we calculate the number of aperture values exceeding the threshold, which represent the number of false positives in the considered annulus. The false positive rate (FPR) is then defined as the ratio of these false positives to the total number of independent realizations. For true positives, we follow Mawet et al. (2014) by first estimating noise levels in the residual cube produced by APCA. We then inject synthetic point sources with fluxes between two and three $\sigma$ of the measured noise. An injected source is considered recovered if its RSM score surpasses the considered detection threshold, and the number of such recoveries within the annulus defines the true positives. The optimal threshold is determined by maximizing the F1 score, ensuring a balance between high true positive recovery and minimal false positives.
Figure 5 shows the evolution of TPR, FPR, and the F1 score for varying thresholds, applied to the cluster 1 center (corresponding to the BD-15 705 dataset) using RSM with APCA at $0\aas@@fstack{\prime\prime}7$ separation. The figure highlights the fact that the optimal F1 score favors minimal FPR and high TPR, thereby effectively distinguishing noise from potential true positives. This metric is highly dataset-specific as it relies heavily on the false positive rate and the true positive rate, the latter also depending on the flux of injected signals, which is influenced by the noise distribution. Consequently, this approach captures the unique characteristics of each dataset and aligns the threshold closely with the specific noise properties of the data. This technique, however, has several limitations. First, while it minimizes the false positive rate, it does not explicitly control the acceptable number of false positives per annulus. Additionally, it assumes all independent realizations are false positives, which may not hold if a real signal is present. In such cases, the maximum F1 score metric tends to exclude the true signal by minimizing false positives. One solution is to remove bright signals by injecting negative fake companions to avoid bias. Finally, this approach requires noise computation, fake companion injections, and a complete RSM processing for each separation, making it computationally expensive.
@@ 5.3 Comparison and reliability of the detection thresholds
Figure 6: Comparison of detection thresholds with false positives and injected companions for the cluster 1 center at an angular separation of $0\aas@@fstack{\prime\prime}6$ and showing the RSM scores of injected fake companions (gold dots), false positives (red crosses), the maximum F1 score threshold (dashed blue line), and the lognormal threshold (solid blue line).
@@
While it arguably provides the most relevant thresholds for a given dataset, the maximum F1 score is computationally demanding to apply on each dataset of the survey, and too specific to generalize across an entire cluster (see more about that in Appendix G). Conversely, the detection thresholds derived from a lognormal distribution are more conservative, computationally efficient, and reflect the noise behavior under similar observing conditions. However, they rely on a predefined distribution, large sample sizes, and do not capture dataset-specific nuances. For the present study, focusing on a large survey, we adopt the threshold derived from the $3\times 10^{-7}$ FAP of the lognormal distribution, and derive it at each separation, for each combination of PSF subtraction technique, and for each cluster. We recommend using the F1 score threshold for smaller samples, while noting its limitations described in Sect. 5.2.
@@ 6 Sensitivity limits
In this study, we adopt the 50% completeness curve described in Dahlqvist et al. (2021b), also known as contrast curves, to assess the sensitivity limits to point-like sources. Such curves indicate the flux at which 50% of injected signals are detected above the selected threshold, here the $3\times 10^{-7}$ FAP from the lognormal distribution, which corresponds to the $5\sigma$ contrast curve when assuming Gaussian noise. Here, we analyze the effects of threshold selection and PSF subtraction technique combinations on contrast curves, define the optimal RSM contrast curve, and evaluate its reliability.
@@ 6.1 Influence of threshold and PSF subtraction techniques
Figure 8: Contrast comparison at 50% completeness using thresholds based on the maximum F1 score (blue), the lognormal RSM noise assumption (orange), and the $5\sigma$ APCA contrast curve (black) across different PSF subtraction techniques applied to the center of cluster 1.
@@
To evaluate how the chosen threshold impacts contrast curves, Fig. 8 presents a comparison for the center of cluster 1. It includes contrast curves derived using the maximum F1 score threshold (blue) and the lognormal threshold (orange) across four combinations of PSF subtraction techniques: APCA, APCA-NMF, APCA-NMF-LOCI, and APCA-NMF-LOCI-FMKLIP-FMLOCI. A standard $5\sigma$ contrast curve based on full-frame PCA with five principal components is shown in black for comparison. The RSM contrast curves show a good level of consistency between the two ways to define the detection threshold. This is especially the case when multi-technique combinations are used (see bottom panels in Fig. 8), while the lognormal threshold leads to slightly worse sensitivity than the F1-score one when RSM is performed with APCA only (Fig. 8 top left) due to the higher threshold in that case. Consistency was also found between the lognormal contrast curves derived here and the contrast curves based on the appearance of a first false positive proposed in Dahlqvist et al. (2022) for the SHARDDS survey (see Appendix E).
@@
Figure 9: Effect of different combinations of PSF subtraction techniques in RSM on the fitted lognormal distributions, detection thresholds, and contrast curves. Top left: Difference of lognormal fit across different PSF subtraction techniques with colors ranging from dark violet for multi combinations to yellow for singular combinations. Top right: Distribution of the $3\times 10^{-7}$ FAP under lognormal distribution for the different combinations across all separations. Bottom left: Contrast curve delivered by the different combinations with 50% completeness.
@@
The choice of which PSF subtraction techniques to combine is an important aspect of the RSM algorithm, which is not straightforward to optimize. Instead of relying on a single combination identified as optimal by the RSM framework, we propose to systematically evaluate the performance of various combinations of PSF subtraction techniques. We analyze the differences in noise distributions in the final frames, the corresponding thresholds, and the contrast curves achievable with each combination. Figure 9 examines these effects across the 28 possible combinations of PSF subtraction techniques, ranging from single PSF subtraction techniques (in yellow) to multi-combinations (in dark violet). Figure 9 (top left) displays the median lognormal fits for each combination, clearly showing narrower noise distributions for multi-combinations and broader distributions for single techniques, indicating improved performance with combined techniques. Figure 9 (top right) presents the thresholds derived from these fits across all combinations and separations. Noise thresholds, on average, range from approximately $10^{-4}$ for multi-combinations to nearly ten for LOCI – exceeding RSM’s maximum score. This renders LOCI non-physical and unreliable when used alone, a consequence of its high false positive rate in RSM for many datasets. Figure 9 (bottom left) illustrates the contrast curves for the various combinations, showing an improvement by a factor of three for multi-combinations. Contrast values with lognormal thresholds exceeding 0.1 in RSM were excluded from the bottom-left panel due to difficulties in achieving convergence. This further demonstrates that incorporating multiple PSF subtraction techniques in RSM significantly enhances detection limits. Moreover, it highlights RSM’s ability to leverage the diverse noise patterns of different PSF subtraction methods, effectively reducing the misclassification of transient speckles as planetary signals. RSM’s efficiency in integrating additional information – either by combining different PSF subtraction techniques or by increasing the number of frames – is evident through narrower noise distributions, improved thresholds, and enhanced detection limits (see also Figure 4(a)).
@@ 6.2 Optimal usage of RSM
Figure 11: Optimal contrast curve achieved by different combinations of PSF subtraction techniques in RSM. Gray curves show the median contrast performance for each combination, computed across all datasets in cluster 1. The red curve highlights the best achievable performance at each separation, given by the median of the minimum contrast values across all combinations. The black curve shows the median of the maximum contrast values delivered by the two best-performing combinations: APCA-NMF-LOCI-FMKLIP and APCA-LOCI-FMKLIP.
@@
Figure 12: Comparison of recovered fake point sources in cluster center 1 using RSM maps with different PSF subtraction technique combinations: (a) APCA-NMF-LOCI-FMKLIP and (b) APCA-LOCI-FMKLIP, compared to the signal-to-noise ratio PCA map with five principal components (c). The RSM maps shown in this figure are thresholded (RSM map minus the separation-dependent threshold). Injected signals recovered above the detection threshold are marked with yellow squares, while unrecovered signals appear in red squares.
Figure 13: Injected fake point sources compared to the optimal RSM contrast curve (black) and PCA five-component contrast curve (red) for cluster center 1